OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find 2f(1) + 3g(4).

OpenStudy (aravindg):

just substitute the values!!

OpenStudy (anonymous):

when you have something like "f(1)", just rewrite the expression using 1 instead of x

OpenStudy (aravindg):

f(1)=1^1+1

OpenStudy (aravindg):

g(4)=3(4)+1

OpenStudy (aravindg):

now cant you do the rest ?

OpenStudy (anonymous):

Subtitute the value to the equation: f (1) = 1^2 + 1 = 1+1=2 g(4) = 3.4 + 1 = 12 + 1 = 13 So, 2f(1) + 3f(4) = 2.2 + 3.13 = ??? (I'm sure you can finish the rest ~_^ )